OpenStudy (anonymous):

find the LCM of 20 and 30... what is the LCM anyways?

OpenStudy (anonymous):

LCM=lowest common multiple

OpenStudy (anonymous):

It is the lowest integer that can be evenly (no decimals) divided by the two given numbers

OpenStudy (anonymous):

10, 5, 1, 30

OpenStudy (anonymous):

2

OpenStudy (lgbasallote):

let me teach you the art of prime decomposition method  so just multiply \[2 \times 5 \times 2 \times 3\] and you get your LCM